# Vertical Semi-Federated Learning for Efficient Online Advertising

### v3

@@ Abstract.
Vertical Federated Learning and Advertising and Knowledge Distillation
copyright: none
@@ 1. Introduction
Vertical Federated Learning (VFL) (vfl_survey1; vfl_survey2) has emerged as a proactive solution (fldssm; blindfl; wei2023fedads; ouyang2024fedud) for leveraging cross-platform user data to enhance recommendation and advertising performance, while preserving privacy and complying with regulations (gdpr). Specifically, in a typical two-party VFL setting(splitNNVP), There is an active party A who holds labels and part of the features and a passive party B who provides additional features. They each privately hold a portion of the model and jointly train a federated model by communicating only hidden features and the corresponding gradients. In this way, Party A can leverage additional data fields from the B-side to improve its own task performance while preserving privacy.
@@
Despite its promise, two main drawbacks hinders its deployment in practical scenarios. (1) Restricted scope on overlapped data: Vanilla VFL trains and serves only overlapped users, who typically constitute a small portion of the population. This limits information richness during training and prevents non-overlapped users from benefiting at all. As a result, the active party has little incentive to deploy a high-cost distributed learning system that yields limited returns, especially when a simpler local model can already serve all users more cost-efficiently. (2) High-cost distributed serving: The distributed nature of VFL introduces extra latency from cross-party data transmission and poses engineering challenges due to heterogeneous network and computational conditions. While such overhead is tolerable during offline training, it becomes nearly infeasible for online inference, where advertising systems demand both high throughput and strict real-time latency (millions of QPS and 10–100 ms per request (scale_rtb_icdm; survey_rtb)). These constraints make VFL deployment in production prohibitively expensive.
@@
These limitations reveal the need for a practical learning framework that can leverage all labeled data while supporting local inference for all users. We term this paradigm Semi-VFL (Vertical Semi-Federated Learning) to highlight its partially federated nature. Since ideal VFL is often impractical to deploy, the goal of Semi-VFL is to outperform vanilla local models. Notably, An effective Semi-VFL solution must address two key challenges: (1) Alleviating field missing: preserving and generalizing passive-party field knowledge from federated training to local inference is essential for maintaining the advantages of VFL; (2) Adapting to the heterogeneous full sample space: effectively combining the imbalanced field information of overlapped and non-overlapped samples is crucial for achieving consistent gains across all users. Several recent studies explore integrating unaligned data or imputing B-side missing fields in the latent space (ouyang2024fedud; kang2022fedcvt), but they mainly aim to improve federated performance rather than support local inference. A few distillation-based approaches (pri_fed; vfedssd) do produce local models compatible with Semi-VFL, yet they rely on direct logit distillation and lacks comprehensive cross-party knowledge transfer.
@@ 3. Experiments
Table 2. The overall results on all sample subsets. Typically, an AUC increase of 0.001 can be considered a significant in CTR Prediction(zhou2018din).
@@ 3.1.1. Datasets
We evaluate our method on two widely used public CTR prediction datasets, Criteo (criteo) and Avazu (criteo). Following the VFL setting in prior work (vfedssd), we partition each dataset’s feature fields into two disjoint subsets to simulate the two parties. Specifically, for Avazu, both parties hold 11 feature fields, while for Fed-Criteo, the active and passive parties have 19 and 20 fields, respectively. We respectively construct 6, 1 and 1 million samples for training validation and testing. Following (kang2022fedcvt), we select earlier half data with older timestamps to construct non-overlapped samples.
@@ 3.1.2. Baselines and Evaluation Setting
We compare our approach against four baselines: (1) Local. A local model is trained solely on the active party’s features, without incorporating any attributes from the passive side. A qualified Semi-VFL method is expected to outperform this baseline. (2) FPD (pri_fed; vfedssd). These pioneering works introduced privileged distillation to enable local inference for all users in the VFL setting. We collectively refer to them as FPD methods. They first train a federated model on shared (aligned) samples and then distill its knowledge into a local model by jointly minimizing the loss over ground-truth labels and soft labels from the federated model. In our implementation, the training objective is defined as $\mathcal{L}_{FPD}=\mathcal{L}^{n}_{CE}+(1-\alpha)\mathcal{L}^{o}_{CE}+\alpha\mathcal{L}^{o}_{KL}$, where $\alpha$ denotes the distillation strength. (3) FedUD* (ouyang2024fedud). FedUD leverages unaligned samples by generating passive-side representations using a pre-trained cross-party feature transfer encoder. Since the original method does not support local inference, we modify it to employ the transfer encoder for both aligned and unaligned samples, enabling local inference in our setting. (4) FedCVT* (kang2022fedcvt). This method leverages unaligned samples during training by imputing B-side hidden representations from aligned samples’ B-side features using its similarity-based mechanism. It effectively learns implicit A-to-B transferable knowledge that can exploit unaligned samples. However, this imputation is not applicable under local inference. Although originally designed for the full VFL scenario, we modify it to use only classifier A’s output, making it a comparable baseline for local inference. Following (wei2023fedads), we implement an efficient, fast version of this method.
We report AUC and logloss results of all baselines evaluated on All sample group under the local inference mode, where only local fields are available during inference.
@@ 3.2.2. Ablation Study
3.2.3. Impact of Non-Overlapped Data Volume
To evaluate the robustness of our methods on different unaligned subsets, we conducted additional experiments on datasets with decreasing proportions of unaligned samples: 70%, 50%, 30%, and 10%. As shown in Figure 2, incorporating more unaligned data generally leads to a performance increase, while the advantage of distillation remains consistent across different data volumes.
Figure 2. Impact of the passive party’s field amount (left) and non-overlapped data volume (right). JPL maintains consistent advantage.
3.2.4. Impact of the Passive Party’s Field Amount
To assess robustness across field configurations, we vary the field sets by using 25%, 50%, and 75% of the fields (Sets A, B, and C) and retrain both teacher and student models. Note that due to differences in field importance and vocabulary distributions, more fields do not necessarily implies better input information. As shown in Figure 2, our method consistently outperforms both Local and FPD across all configurations. Since the Local model is unaffected by this operation, its performance bars remain identical.
@@ 3.2.5. Effectiveness across Different Architectures
To evaluate the generalizability of our framework to different backbone architectures, we design two variants of Local, FPD and JPL based on DeepFM (guo2017deepfm) and DCN (wang2017DCN). These architectures are applied to the student’s local bottom model, and new students are trained while keeping the teacher model unchanged. As shown in Table 4, JPL consistently demonstrates superior performance, validating the effectiveness of our framework under heterogeneous teacher–student setups and across various local model architectures.

### v4

@@ Abstract.
Advertising; Vertical Federated Learning; Knowledge Distillation
copyright: acmcopyrightjournalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792940isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Online advertising
@@ 1. Introduction
Vertical Federated Learning (VFL) (Liu et al., 2024) has emerged as a proactive solution (Wei et al., 2023; Ouyang et al., 2024; Li et al., 2024) for leveraging cross-platform user data to enhance recommendation and advertising performance, while preserving privacy and complying with regulations. Specifically, in a typical two-party VFL setting(Ceballos et al., 2020), There is an active party A who holds labels and part of the features and a passive party B who provides additional features. They each privately hold a portion of the model and jointly train a federated model by communicating only hidden features and the corresponding gradients. In this way, Party A can leverage additional data fields from the B-side to improve its own task performance while preserving privacy.
@@
Despite its promise, two main drawbacks hinders its deployment in practical scenarios. (1) Restricted scope on overlapped data: Vanilla VFL trains and serves only overlapped users, who typically constitute a small portion of the population. This limits information richness during training and prevents non-overlapped users from benefiting at all. As a result, the active party has little incentive to deploy a high-cost distributed learning system that yields limited returns, especially when a simpler local model can already serve all users more cost-efficiently. (2) High-cost distributed serving: The distributed nature of VFL introduces extra latency from cross-party data transmission and poses engineering challenges due to heterogeneous network and computational conditions. While such overhead is tolerable during offline training, it becomes nearly infeasible for online inference, where advertising systems demand both high throughput and strict real-time latency (millions of QPS and 10–100 ms per request (Shen et al., 2015)). These constraints make VFL deployment in production prohibitively expensive.
@@
These limitations reveal the need for a practical learning framework that can leverage all labeled data while supporting local inference for all users. We term this paradigm Semi-VFL (Vertical Semi-Federated Learning) to highlight its partially federated nature. Since ideal VFL is often impractical to deploy, the goal of Semi-VFL is to outperform vanilla local models. Notably, An effective Semi-VFL solution must address two key challenges: (1) Alleviating field missing: preserving and generalizing passive-party field knowledge from federated training to local inference is essential for maintaining the advantages of VFL; (2) Adapting to the heterogeneous full sample space: effectively combining the imbalanced field information of overlapped and non-overlapped samples is crucial for achieving consistent gains across all users. Several recent studies explore integrating unaligned data or imputing B-side missing fields in the latent space (Ouyang et al., 2024; Kang et al., 2022), but they mainly aim to improve federated performance rather than support local inference. A few distillation-based approaches (Ren et al., 2022; Li et al., 2022) do produce local models compatible with Semi-VFL, yet they rely on direct logit distillation and lacks comprehensive cross-party knowledge transfer.
@@ 3. Experiments
Table 2. The overall results on all sample subsets. Typically, an AUC increase of 0.001 can be considered a significant in CTR Prediction(Zhou et al., 2018).
@@ 3.1.1. Datasets
We evaluate our method on two widely used public CTR prediction datasets, Criteo and Avazu under a VFL data partition setting as in  (Li et al., 2022). Specifically, for Avazu, both parties hold 11 feature fields, while for Criteo, the active and passive parties have 19 and 20 fields, respectively. We respectively construct 6, 1 and 1 million samples for training validation and testing. Following (Kang et al., 2022), we select earlier half data to construct non-overlapped samples.
@@ 3.1.2. Baselines and Evaluation Setting
We compare our approach against four baselines: (1) Local. A local model is trained solely on the active party’s features. A qualified Semi-VFL method is expected to outperform this baseline. (2) FPD (Ren et al., 2022; Li et al., 2022). These pioneering works first train a federated model on overlapped samples and then distill logit knowledge into a local model, thus we collectively refer to them as FPD (federated privileged distillation) methods. In our implementation, the training objective is defined as $\mathcal{L}_{FPD}=\mathcal{L}^{n}_{CE}+(1-\alpha)\mathcal{L}^{o}_{CE}+\alpha\mathcal{L}^{o}_{KL}$, where $\alpha$ denotes the distillation strength. (3) FedUD* (Ouyang et al., 2024). FedUD leverages unaligned samples by generating passive-side representations using a pre-trained cross-party feature transfer encoder. Since the original method does not support local inference, we modify it to employ the transfer encoder for both aligned and unaligned samples. (4) FedCVT* (Kang et al., 2022). This method leverages unaligned samples during training by imputing B-side hidden representations from aligned samples’ B-side features using its similarity-based mechanism. It effectively learns implicit A-to-B transferable knowledge that can exploit unaligned samples. However, this imputation is not applicable under local inference. Although originally designed for the full VFL scenario, we modify it to use only classifier A’s output, making it a comparable baseline for local inference. Following (Wei et al., 2023), we implement an efficient, fast version of this method. We report AUC and logloss results of all baselines evaluated on All sample group under the local inference mode, where only local fields are available during inference.
@@ 3.2.2. Ablation Study
3.2.3. Impact of Data Volume and Field Amount
As shown in Figure 2(left), we conducted experiments on datasets with decreasing proportions of unaligned samples: 70%, 50%, 30%, and 10%. It shows that incorporating more unaligned data generally leads to a performance increase, while the advantage of distillation remains consistent across different data volumes. While in Figure 2(right), we vary the field sets by using 25%, 50%, and 75% of the fields (Sets A, B, and C) to assess robustness across field configurations. Note that due to differences in field importance and vocabulary distributions, more fields do not necessarily implies better input information. The results show that our method consistently outperforms both Local and FPD across all configurations. Since the Local model is unaffected by this operation, its performance bars remain identical.
Figure 2. Impact of non-overlapped data volume (left) and the passive party’s field set (right). JPL maintains consistent advantage.
@@ 3.2.4. Effectiveness across Different Architectures
To evaluate the generalizability of our framework to different backbone architectures, we design two variants of Local, FPD and JPL based on DeepFM (Guo et al., 2017) and DCN (Wang et al., 2017). These architectures are applied to the student’s local bottom model, and new students are trained while keeping the teacher model unchanged. As shown in Table 4, JPL consistently demonstrates superior performance, validating the effectiveness of our framework under heterogeneous teacher–student setups and across various local model architectures.
@@ 4. Conclusion
Acknowledgements.
This work is supported in part by the National Natural Science Foundation of China under Grant 62571298. Thanks to Tencent Rhino-bird Elite Talent Program for the early start of this project.